Consider the following hypotheses: H0: p ? 0.38 HA: p < 0.38

Compute the p-value based on the following sample information

a. x = 22; n = 74

b. x = 110; n = 300

c. pbar = 0.34; n = 50

d. pbar = 0.34; n = 400
Answer #1

Consider the following hypotheses: H0: p ? 0.38 HA: p < 0.38

To find the p-values for following cases:

a) x = 22; n = 74

\hat p =\frac{x}{n}=\frac {22}{74}=0.2973

Since given test is left tailed test,

Pvalue=P(\hat p < 0.2973)

              =P(\frac{\hat p - p}{\sqrt {\frac {p(1-p)}{n}}}< \frac{0.2973 - 0.38}{\sqrt {\frac {0.38(1-0.38)}{74}}})

              =P(Z < -1.47)

              =1-P(Z < 1.47)

              =1-0.9292

              =0.0708

Hence, P-value = 0.0708

                  

b) x = 110; n = 300

\hat p =\frac{x}{n}=\frac {110}{300}=0.3667

Since given test is left tailed test,

Pvalue=P(\hat p < 0.3667)

              =P(\frac{\hat p - p}{\sqrt {\frac {p(1-p)}{n}}}< \frac{0.3667 - 0.38}{\sqrt {\frac {0.38(1-0.38)}{300}}})

              =P(Z < -0.47)

              =1-P(Z < 0.47)

              =1-0.6808

              =0.3192

Hence, P-value = 0.3192

c) pbar = 0.34; n = 50

Since given test is left tailed test,

Pvalue=P(\hat p < 0.34)

              =P(\frac{\hat p - p}{\sqrt {\frac {p(1-p)}{n}}}< \frac{0.34 - 0.38}{\sqrt {\frac {0.38(1-0.38)}{50}}})

              =P(Z < -0.58)

              =1-P(Z < 0.58)

              =1-0.7190

              =0.2810

Hence, P-value = 0.2810

d) pbar = 0.34; n = 400

Since given test is left tailed test,

Pvalue=P(\hat p < 0.34)

              =P(\frac{\hat p - p}{\sqrt {\frac {p(1-p)}{n}}}< \frac{0.34 - 0.38}{\sqrt {\frac {0.38(1-0.38)}{400}}})

              =P(Z < -1.65)

              =1-P(Z < 1.65)

              =1-0.9505

              =0.0495

Hence, P-value = 0.0495